Accountants at the firm Michael​ Vest, Chartered​ Accountants, believed that several traveling executives were submitting unusually high travel vouchers when they returned from business trips.​ First, they took a sample of 200 vouchers submitted from the past year. Then they developed the following​ multiple-regression equation relating expected travel cost to number of days on the road ​(x 1​) and distance travelled ​(x 2​) in​ kilometres: ModifyingAbove y with caretequals​$85.00plus​$46.50x 1plus​$0.45x 2. The coefficient of correlation computed was 0.68. ​a) If Wanda Fennell returns from a 320​-kilometres trip that took her out of town for 6 ​days, the expected amount that she should claim as expensesequals​$ 256.9 ​(round your response to two decimal​ places). ​b) Fennell submitted a reimbursement request for ​$785. What should the accountant​ do? After the 320​-kilometres and 6​-day ​trip, Fennell submitted a reimbursement request for ​$785​, which is different than the predicted value. The accountant should not question question not question the voucher as the request is within expectations within expectations much higher than expected . ​c) Should any other variables be​ included? Which​ ones? Why? Additional variables that could be included in the model are ​(select the choice that has the factor you would​ consider): A. Type of travel​ (air/car) B. Dress code​ (formal/business casual) C. Gender of the executive​ (male/female) The percentage of variation in the trip cost that can be explained by the modelequals 0​% ​(round your response to two decimal ​places)​, which suggests that the model is not particularly a good one not particularly a good one particularly a good one .
Answer #1

(a)

Claim (Y) = 85 + 46.5 x1 + 0.45 x2

Given, x1 = 6 days and x2 = 320 Km

So, Y = 85 + 46.5*6 + 0.45*320 = $508

(b)

The accountant should question the voucher as the request amount of 785 is higher than the prediction which is 508

(c)

Other relevant variables can be:

  • A. Type of travel​ (air/car)

(d)

The percentage of variation in the trip cost that can be explained by the model = R2 = 0.68^2 = 46.24%.

This suggests that the model is not particularly a good one as the value is even less than 50%.